OpenStudy (bruised1998):

5^x+8=1008 round to the nearest ten-thousandth?

Vocaloid (vocaloid):

subtract 8 from both sides to get 5^x = 1000 then apply a logarithm to both sides log_5(5^x) = log_5(1000)

Vocaloid (vocaloid):

x = log(1000)/log5 which you can find on a calculator

OpenStudy (cal.lavender):

8 + 5^x = 1008 5^x = 1000 apply log on both sides x log 5 = log 1000 x = 3 / log 5 x = 3 / 0.6989 x = 4.2920

OpenStudy (bruised1998):

thank you !!

OpenStudy (cal.lavender):

np c: